Build a factor tree to get the prime factorization of each number.
1. 280
2. 145
3. 117
4. 159
5. 134

Answer by tommyt3rd(5050)

diagrams are really hard to do on this software...so I'll do the first (only) like this:

280 = 2*140=
2*2*70 =
2*2*2*35=
2*2*2*5*7

:)